Sol Summary – November 21st

Crew 286 Sol 08 Summary Report 20-NOV-2023

Sol: 08

Summary Title: “A Fine Day Sampling Mars.”

Author’s name: Roger Gilbertson

Mission Status: Nominal

Sol Activity Summary: This morning’s EVA #8 to collect samples from three locations went smoothly. We prepared efficiently and exited the airlock at the planned time, driving out of the campus and south on Cow Dung Road. Gee used the iPhone GPS and directed us to stop on the road near the coordinates provided by the previous researchers. We trekked eastward, and located each sampling position with relative ease. The locations where the first two samples were previously collected appeared clearly on the ground, having only received minor erosion and filling in since their April sampling. The third sample, been in a stream bed, were not apparent, but we took samples at the indicated coordinates. At each location we collected and labeled samples, and photographed each before and after digging.

We returned to the vehicles with an hour remaining, and asked for a time extension so that filmmaker Hugo could continue to film content for his documentary, including shots taken by him standing at the roadside and filming as we drove past in the rovers. We continued northwards until we filmed rovers entering the campus, and astronauts entering the airlock.

We have cleaned both floors of the mud which entered after the rains. Science sampling continued, the purple bacteria continues to show steady growth. The second half of the metal casting mold was dried and trimmed. Then both halves were joined, and the first bismuth casting performed. It was removed from the mold successfully, and a second casting was performed prior to this report. Time permitting, a third cast will be made this evening, and then the three pieces can be assembled tomorrow into the final structure. Drone test flights were successfully performed in the tunnel between the habitat and RAM, in preparation for tomorrow’s EVA.

SOL 9 PLAN: Morning EVA #9, Afternoon EVA #10, completion of science and metal casting projects, begin drafting our Final Mission Report.

SOL 10 PLAN (Thursday, last full day of sim): The science program will continue to monitor bacterial growth. Plans have been made for our final sols, including our Thanksgiving dinner. Our EVA 10 proposal is in development and proposes three crew members driving the MASH vehicle southwards, parking at the Barrainca Butte road, then conducting a 2-person walking EVA in the vicinity with CapCom operating in the MASH. The EVA walkers will conduct comms test from nearby hilltops to hopefully reach the stations via the CH1 repeater, then return to the MASH vehicle. After a rest and recharge, a second brief demonstration EVA may be conducted before returning to the habitat. We estimate the entire excursion to be about 4 hours from exit to return. Then we will complete our Final Mission Report, and celebrate with a martian Thanksgiving dinner.

EXIT PLAN: At present we plan to exit sim after waking Friday morning. We will clean, inventory, pack, and load the Crew Car. Then, as time allows, we would like to borrow the Habitat Car and conduct an out-of-sim drive to the Special Area to the north. We will return to campus by 4 PM, then depart in the Crew Car for Grand Junction.

Anomalies in work: none
Weather: sunny

Crew Physical Status: nominal

EVA: EVA 8 completed, request for EVA 9 and EVA 10 submitted, EVA 11 outlined above

Reports to be filed: Sol Summary, Journalist report, Photos, Operations report, Green Hab report, EVA 8 report, EVA 9 and 10 requests.

Support Requested: none

Copyright © The Mars Society. All rights reserved. | Main Site